Adult tickets for a play cost $20 and child tickets cost $13. If there were 29 people at a performance and the theater collected $510 from ticket sales, how many children attended the play?
Answer by checkley71(8405)

A+C=29 A=29-C
20A+13C=510 NOW SUBSTITUTE (29-C) FOR A IN THIS EQUATION 7 SOLVE FOR C
20(29-C)+13C=510
580-20C+13C=510
-7C=510-580
-7C=-70
C=-70/-7
C=10 CHILDREN TICKETS SOLD.
29-10=19 ADULT TICKETS SOLD.
PROOF
20*19=10*13=510
280+130=510
510=510